# How Good is ChatGPT in Giving Advice on Your Visualization Design?

### v3

@@ Abstract.
Data visualization creators often lack formal training, resulting in a knowledge gap in design practice. Large-language models like ChatGPT, with their vast internet-scale training data, offer transformative potential in addressing this gap. In this work, we first utilized the VisGuide forum, a repository of data visualization questions, to compare ChatGPT-generated responses with Human replies. We additionally conducted a user study delving into practitioners’ reactions and attitudes toward ChatGPT as a visualization assistant. Participants, who brought their visualizations and questions, received feedback from both Human experts and ChatGPT in a randomized order. The results highlight the unique advantages and disadvantages of ChatGPT, such as its ability to quickly provide a wide range of design options based on a broad knowledge base, while also revealing its limitations in terms of contextual understanding, interaction dynamics, and perceived trustworthiness of responses. From these insights, we derive design considerations for LLM-based design feedback systems.
@@ 1. Introduction
However, open questions remain regarding the extent of LLMs’ current knowledge of visualization best practices, the quality and practical applicability of their suggestions, and how these suggestions differ from those provided by human expert visualization designers. Addressing these open questions, we have formulated two specific research questions to guide our investigation:
RQ1. Can ChatGPT rival Human expertise in data visualization knowledge?
@@
RQ2. How would visualization practitioners perceive ChatGPT’s design feedback?
@@
We employ a mixed-methods approach (Fig. 1) to explore the potential of ChatGPT in offering data visualization design knowledge. We selected ChatGPT as a representative proxy for LLMs due to its widespread use and generally superior performance compared to other LLMs (Borji and Mohammadian, 2023; Ahmed et al., 2023; Giannakopoulos et al., 2023; Lim et al., 2023). First, we investigated posts and discussions at the VisGuides forum (Diehl et al., 2018), which captures public questions and responses about visualization design. We asked the questions to ChatGPT and compared its responses to the Human responses on VisGuides across six key metrics: coverage, topicality, breadth, clarity, depth, and actionability. We found that ChatGPT’s performance is comparable to, and often better than, Human responses, especially in providing clear and comprehensive answers that cover all questions asked. In contrast, human responses typically offered specific suggestions, like citing references beyond basic improvements, but showed greater variability due to multiple respondents involved.
@@
While the VisGuide analysis provided insight into ChatGPT’s response quality (RQ1), it fell short of capturing practitioners’ reactions, attitudes, and overall experiences (RQ2). Thus, we conducted a qualitative user study wherein we invited practitioners to partake in a comparative assessment of Human expert feedback and ChatGPT-generated feedback. For this study, participants were invited to bring their own visualization designs and projects to our sessions for questions and feedback. Interestingly, all participants preferred discussing with Human experts, valuing their ability to engage in fluid conversations and provide tailored recommendations. However, participants appreciated the broad knowledge of ChatGPT, particularly its ability for quick brainstorming and ideation from a neutral standpoint, while acknowledging its limitations in text-based communication.
@@
Both studies consistently shed light on the strengths and weaknesses of ChatGPT and Humans. ChatGPT excelled at providing broad suggestions, whereas Humans demonstrated a better grasp of nuanced details for more in-depth ideas. A stronger preference for Humans in the feedback study was linked to the more consistent opinions and natural conversational dynamics of Human experts. These findings also point to areas for future enhancement of LLMs, including enhancing visual perception and proactive discussion capabilities to better tackle the nuanced contexts of design problems. Furthermore, we discuss the potential role of an LLM-based knowledge companion in visualization practice and, more broadly, in education.
@@
An evaluation framework for assessing the knowledge of LLMs in data visualization design, based on metrics derived from the analysis of the VisGuides forum.
Insights into LLM-based chatbots in visualization practice via the lens of ChatGPT analysis to highlight their strengths and weaknesses compared to Human counterparts.
Future considerations for enhancing the user interface of LLM-based chatbots as design companions, aiming for improved human-AI interaction.
@@ 2.2. Evaluating LLMs’ Knowledge Capacity
A number of benchmark tasks have been proposed to assess the general knowledge and reasoning capabilities of language models in a more holistic manner (Liang et al., 2022; Chang et al., 2023). These include tasks that require models to answer broad-coverage factoid questions based on pre-existing knowledge (Liang et al., 2022). Other benchmarks evaluate numeric, analogical, and logical reasoning skills across different modalities (Qin et al., 2023; Zhang et al., 2023b), while some others examine the ability of LLMs to generalize knowledge to new concepts (Wei et al., 2022a). Assessing models on such diverse reasoning skills can provide deeper insight into their adaptability, knowledge gaps, and limitations. In addition to assessing such capabilities, an important area of research has focused on evaluating potential harms and biases in large language models. Some studies have performed targeted adversarial evaluations to expose problematic biases encoded in the models’ training data (Nadeem et al., 2021; Tan and Celis, 2019; Gehman et al., 2020).
@@
Recent studies have initiated the exploration of Large Language Models’ (LLMs) applicability in the realm of data visualization. Some investigations have examined their capacity to generate data visualizations (Paik, 2023; Hutchinson et al., 2023), facilitate exploratory data analysis (Dibia, 2023), and create stylistic data graphics (Dibia, 2023), in addition to generating graph drawings (Bartolomeo et al., 2023). Furthermore, studies have delved into the usability of LLM-based code generation in coding data visualizations (Vaithilingam et al., 2022), as well as understanding visualization practitioners’ perceptions regarding generative text-to-image models (Schetinger et al., 2023).
Unlike the previous work, we focus on evaluating the design knowledge capacity of LLMs. This facet is one of the primary pillars of data visualization skills, extending beyond the mere creation of data visualizations. In other words, this work delves into the cognitive aspects of LLMs, assessing not just their visualization code output but the reasoning and principles underlying their suggestions.
@@ 3.1. Data Collection
After compiling the list of qualifying questions, we presented each query to ChatGPT in May 2023, accompanied by a role-playing prompt designed as follows:
@@
To streamline our analysis, we created a spreadsheet where each row included a question, the corresponding ChatGPT response (GPT 3.5), and a compilation of Human responses. Each question had 1.54 Human responses on average ( = 1.03). The average approximate word count of the Human responses was 235.35 ( = 289.62) with a maximum of 1721 words and minimum of 13 words. The average approximate word count of the ChatGPT responses was 463.42 ( = 189.34) with a maximum of 1192 words and minimum of 115 words.
@@ 3.2. Analysis Decisions
We selected ChatGPT as a representative of LLM-based chatbots for several reasons. While these chatbots use the same transformer architectures, ChatGPT is the most widely adopted in practice. Also, in our early pilot experimentation with Bard (now Gemini), Claude, and Bing (which uses the same GPT model as ChatGPT), ChatGPT generally provided responses that were better aligned in terms of style and content. For example, Bing often veered off by drawing information from various web sources, which were frequently inferior to ChatGPT’s internal knowledge responses. Furthermore, Bard’s recommendations were narrower in scope and often not applicable, like suggesting the addition of a trend line to a Gantt Chart. Overall, our impression was that these systems are mostly comparable, with ChatGPT having a slight edge over the others. Several studies also indicated that ChatGPT demonstrates superior performance compared to other LLMs in terms of a variety of task questions involving reasoning, facts, and fairness (Borji and Mohammadian, 2023; Ahmed et al., 2023). That said, we plan to establish a comparative benchmark which we discuss in the future work section.
@@
3.3. Analysis Method
Two researchers performed an open coding process on a randomly selected 20% sample of the questions and the corresponding ChatGPT and Human responses. This process resulted in the development of six evaluation metrics: breadth, clarity, depth, actionability, coverage, and topicality.
These metrics were assessed using a Likert scale that ranged from 1 to 5. For instance, a rating of 1 indicated narrow responses, while a rating of 5 denoted very broad responses in the context of the breadth metric. The researchers initially assigned scores to the 20% sample, engaging in discussions to resolve any discrepancies and establish consensus definitions and criteria for each score within every metric to ensure consistency. Following this, one of the researchers proceeded to score the remaining question responses. Throughout this coding process, they held continuous meetings to discuss any new issues and questions to ensure consistency and quality of codes and scores, refining metric definitions and score applicability.
Given that some questions elicited multiple Human responses while others received only one, We created a composite score to evaluate the collective quality of the Human responses. Furthermore, any elements of the Human responses that extended beyond the capabilities of ChatGPT v3.5, such as providing citations to related literature, links to external resources, visual examples, or posing follow-up questions, were recorded. We also categorized whether each question sought feedback on the questioner’s visualizations or inquired about general visualization design knowledge.
@@ 3.4.1. Ratings on Quality Metrics
Both ChatGPT and Humans excelled in covering the entirety of the question, with ChatGPT slightly outperforming Humans. ChatGPT scored 4.92 ( = 0.33), while Humans scored 4.37 ( = 0.95). The lower score often stemmed from the Human respondent’s failure to address all facets of the question. For instance, in Figure 3C, the user raised a two-part question, “…would you recommend all the states be placed on the Y-axis instead of grouped as regions? If so, which visual layouts would be effective with dealing with the large number of states?…”. However, the Human response only addressed the second part of the question, suggesting a heat map as a useful visual layout.
@@
This score indicates the degree to which the response stays on topic. ChatGPT averaged 4.87 ( = 0.45), while Humans averaged 4.72 ( = 0.59). There were moments when both ChatGPT and Humans strayed from the main topic. For instance, when a user inquired about the application of visual variables in VR visualizations,111visguides:visual-variables-in-vr ChatGPT not only addressed this question but also came up with a hypothetical question asking what factors to consider when choosing such visual variables. Human respondents also occasionally veered off-topic by offering unsolicited advice and extending their responses beyond the original question’s scope.
@@
Breadth measures how widely a response explores various ideas, concepts, options, or perspectives. In this category, ChatGPT outperformed Human responses by a significant margin. ChatGPT achieved an average score of 4.62 ( = 0.61), while Human responses averaged only 3.35 ( = 1.13). For example, in Figure 3A, when the user inquired about color palette guidelines, ChatGPT provided an extensive list of potential approaches, encompassing data type, color harmony, accessibility, etc. In contrast, the Human respondent to this question concentrated primarily on the creation of domain-specific color maps, offering in-depth insights into the process.
@@
The clarity score assesses how easily a reader can comprehend a response, taking into account conciseness, lack of verbosity, and the organization of the response structure. Notably, ChatGPT consistently presented its responses in a well-structured list. This contributed to ChatGPT receiving a high clarity score, averaging 4.93 ( = 0.28). Human responses also demonstrated good clarity with an average score of 4.30 ( = 0.79). They received lower ratings when the responses were overly brief, making it difficult to discern the specific reference points.
Depth
The depth score evaluates the extent of explanations, expertise, and valuable insights in a response. Human responses averaged a score of 3.52 ( = 1.07), while ChatGPT’s average was 3.44 ( = 0.71). Although the scores are comparable to each other, a distinct difference in depth between a ChatGPT response and a Human response is evident in Figure 3B. The ChatGPT response primarily focuses on general, surface-level improvements applicable to any visualization, such as adding labels and legends. In contrast, the Human response often provides more tailored and specific suggestions for enhancing the visualization, as demonstrated by the recommendation to “put a label before the numerical attribute like ’Estimated Production: 1000’.” This detailed guidance directly pertains to the unique aspects of the visualization in question. However, Human responses also displayed greater variability. For instance, when asked about color clarity and alternative data representation to reduce clutter, a Human response simply confirmed the current design choices are fine222visguides:latitude-distribution-of-solar-and-wind-farms-uk. The lower scores of such poor quality single responses pulled the average down.
@@
The actionability score assesses whether the response offers guidance that can be readily implemented in the visualization. This metric is not applicable to visualization guideline questions. The actionability scores were quite similar: Human responses had an average score of 3.86 ( = 0.94), while ChatGPT achieved an average of 3.99 ( = 0.61). Overall, ChatGPT received a high score for actionability, as it consistently provided a list of recommendations. However, the score was occasionally reduced due to the generic or vague nature of these suggestions. For instance, in Figure 3B, ChatGPT offered advice for addressing label clutter, selecting colors, and refining the legend. Nevertheless, these recommendations were explained in a manner that made them applicable to most visualizations. The actionability scores for Human responses tend to be low when the Human refrains from offering recommendations and instead validates all of the user’s choices333visguides:comparison-of-acute-hepatitis-from-tycho-dataset.

### v4

@@ Abstract.
Data visualization creators often lack formal training, resulting in a knowledge gap in design practice. Large-language models such as ChatGPT, with their vast internet-scale training data, offer transformative potential to address this gap. In this study, we used both qualitative and quantitative methods to investigate how well ChatGPT can address visualization design questions. First, we quantitatively compared the ChatGPT-generated responses with anonymous online Human replies to data visualization questions on the VisGuides user forum. Next, we conducted a qualitative user study examining the reactions and attitudes of practitioners toward ChatGPT as a visualization design assistant. Participants were asked to bring their visualizations and design questions and received feedback from both Human experts and ChatGPT in randomized order. Our findings from both studies underscore ChatGPT’s strengths—particularly its ability to rapidly generate diverse design options—while also highlighting areas for improvement, such as nuanced contextual understanding and fluid interaction dynamics beyond the chat interface. Drawing on these insights, we discuss design considerations for future LLM-based design feedback systems.
@@ 1. Introduction
However, open questions remain regarding the extent of LLMs’ current knowledge of visualization best practices, the quality and practical applicability of their suggestions, and how these suggestions differ from those provided by human counterparts. Addressing these open questions, we have formulated two following research questions to guide our investigation:
RQ1 (Quantitative evaluation): How well does ChatGPT respond to visualization design questions posed in online communities compared to practitioners in the wild?
@@
RQ2 (Qualitative experience): How do visualization practitioners perceive ChatGPT’s design feedback compared to those offered by seasoned professionals?
We employ a mixed-methods approach (Fig. 1) to tackle these questions, capturing both breadth (RQ1) and depth (RQ2). RQ1 focuses on a more systematic evaluation using a broader and more realistic knowledge baseline, while RQ2 centers on the qualitative experience of receiving feedback from ChatGPT, assessed against the high standards of domain experts. In our comparisons, seasoned professionals represent trusted colleagues who bring domain expertise and rigor, whereas online practitioners provide a practical, real-world perspective, recognizing that not everyone has access to expert advice (Choi et al., 2023b). We selected ChatGPT as a representative proxy for LLMs due to its widespread use and generally superior performance compared to other LLMs (Borji and Mohammadian, 2023; Ahmed et al., 2023; Giannakopoulos et al., 2023; Lim et al., 2023).
@@
For the first question, we investigated posts and discussions at the VisGuides forum (Diehl et al., 2018), which captures public questions and responses about visualization design. We fed the questions to ChatGPT and compared its responses to the Human responses on VisGuides across six key metrics: coverage, topicality, breadth, clarity, depth, and actionability. We found that ChatGPT’s response is comparable to, and often better than, Human responses, especially in providing clear and comprehensive answers that cover all questions asked. In contrast, human responses typically offered specific suggestions, like citing references beyond basic improvements, but showed greater variability due to multiple respondents with diverse backgrounds involved. Given our analysis pertains to specific versions of ChatGPT, we discuss how we can devise an evaluation pipeline that uses LLMs for automatic response scoring.
@@
For the second question, we conducted a qualitative user study in which we invited practitioners to participate in a comparative evaluation of Human expert feedback and ChatGPT-generated feedback (GPT-3.5 Turbo). For this study, participants were invited to bring their own visualization designs and projects to our sessions for questions and feedback. Most study participants preferred discussing with Human experts, valuing their ability to engage in fluid conversations and provide tailored recommendations. However, participants also appreciated the broad knowledge of ChatGPT, particularly its ability for quick brainstorming and ideation from a neutral standpoint, while acknowledging its limitations in deeper contextual understanding and its lack of critical feedback.
@@
Both studies consistently shed light on the strengths and weaknesses of ChatGPT and Humans. ChatGPT excelled at providing broad suggestions, whereas Humans demonstrated a better grasp of nuanced details for more in-depth ideas. A stronger preference for Humans in the feedback study was also linked to the more consistent opinions and fluid interaction dynamics of Human experts. These findings also point to areas for future enhancement of LLMs, such as supporting advanced prompting with better visual perception and fostering proactive and critical discussion capabilities to more effectively handle the intricate contexts of design challenges. We distill these lessons into design considerations for future design feedback systems using LLMs. Furthermore, we discuss the potential role of an LLM-based design knowledge companion in visualization education.
@@
Evaluation framework for assessing the capabilities of LLMs in providing data visualization design assistance, incorporating both qualitative and quantitative metrics.
Key insights into the role of LLM-based chatbots, exemplified by ChatGPT, in visualization practice, highlighting their strengths and limitations compared to Human counterparts.
Design considerations for improving the user interface of LLM-based chatbots as collaborative design companions, fostering more effective human-AI interaction.
@@ 2.2. Evaluating LLMs’ Knowledge Capacity
A number of benchmark tasks have been proposed to assess the general knowledge and reasoning capabilities of language models in a more holistic manner (Liang et al., 2022; Chang et al., 2023). These include tasks that require models to answer broad-coverage factoid questions based on pre-existing knowledge (Liang et al., 2022). Other benchmarks evaluate numeric, analogical, and logical reasoning skills across different modalities (Qin et al., 2023; Zhang et al., 2023), while some others examine the ability of LLMs to generalize knowledge to new concepts (Wei et al., 2022a). Assessing models on such diverse reasoning skills can provide deeper insight into their adaptability, knowledge gaps, and limitations. In addition to assessing such capabilities, an important area of research has focused on evaluating potential harms and biases in large language models. Some studies have performed targeted adversarial evaluations to expose problematic biases encoded in the models’ training data (Nadeem et al., 2021; Tan and Celis, 2019; Gehman et al., 2020).
@@
Many recent studies have explored the applicability of LLMs in the realm of data visualization. Inala et al. examine the potential impact of generative AI on the data analysis process, encompassing data collection, transformation, and exploration (Inala et al., 2024). DataFormulator 1 & 2 focus on facilitating data transformation workflows for creating data visualizations (Wang et al., 2023c, 2024c). On the other hand, several tools leverage LLMs to provide tailored natural language interfaces for generating data visualizations, including LIDA (Dibia, 2023), DynaVis (Vaithilingam et al., 2024), and ChartGPT (Tian et al., 2024). Other studies have evaluated LLMs’ capabilities in eliciting design preferences (Wang et al., 2024a), assisting in chart interpretation (Wang et al., 2024b), detecting misleading charts (Lo and Qu, 2024; Alexander et al., 2024), performing visualization literacy tasks (Bendeck and Stasko, 2024; Choe et al., 2024), generating visualization code (Chen et al., 2024), and troubleshooting visualization generation code (Shen et al., 2024). Furthermore, user studies have examined how ChatGPT is utilized in classrooms (Kim et al., 2024), visualization practitioners’ perceptions of text-to-image models (Schetinger et al., 2023), the impact of AI assistance on analysis workflows (Gu et al., 2024), and its effects on visualization coding performance (Vaithilingam et al., 2022).
The two pillars of data visualization skills are design and implementation. Most previous studies have focused on data visualization generation and implementation. In contrast, this work focuses on evaluating the design feedback capabilities of LLMs. Similar efforts have recently been explored in user interface design, where researchers have leveraged existing guidelines (Duan et al., 2024b) and critique datasets (Duan et al., 2024a) to perform design evaluations. In our study, we use real-world design questions from practitioners to conduct evaluations.
@@ 3.1. Data Collection
After compiling the list of qualifying questions, we fed each query to ChatGPT 3.5 in May 2023 and with images to ChatGPT 4 with Vision in May 2024 (accompanied by images), accompanied by a role-playing prompt designed as follows:
@@
Please format your reply as JSON objects with keys: Question and Response.
Question: xxx
Response: xxx
There may be more than one question, format each as a new object.
To streamline our analysis, we created a spreadsheet where each row included a question, the corresponding ChatGPT responses (3.5 and 4V) and a compilation of Human responses. Each question had 1.54 Human responses on average ( = 1.03). The average approximate word count of the Human responses was 235.35 ( = 289.62) with a maximum of 1721 words and minimum of 13 words. The average approximate word count of the ChatGPT 3.5 responses was 463.42 ( = 189.34) with a maximum of 1192 words and minimum of 115 words. For ChatGPT 4, the responses averaged around 255.03 words ( = 88.58), with lengths spanning from 110 to 491 words.
3.2. Analysis Method
Two researchers performed an open coding process on a randomly selected 20% sample of the questions and the corresponding ChatGPT and Human responses. They iteratively agreed on and refined six evaluation metrics: coverage, topicality, breadth, depth, clarity, and actionability. Since design is a complex problem-solving task, we believe that qualitative human-centered metrics are more appropriate for capturing the holistic quality of responses compared to traditional utility metrics such as accuracy. These metrics were assessed using a Likert scale that ranged from 1: Very poor to 5: Very good. Table 1 presents our codebook of metric definitions and scoring rubrics.
| Metric & Definition | Score & Description |
| Coverage assesses whether the response addresses all questions posed by the user. | 1: Addresses 0–20% of the user’s questions. |
| 2: Addresses 20–40% of the user’s questions. | |
| 3: Addresses 40–60% of the user’s questions. | |
| 4: Addresses 60–80% of the user’s questions. | |
| 5: Addresses 80–100% of the user’s questions. | |
| Topicality indicates how closely the response stays on topic. | 1: Mostly or entirely off-topic. |
| 2: Partially addresses the topic but frequently digresses. | |
| 3: Somewhat relevant; minor deviations from the topic. | |
| 4: Mostly on-topic; occasional tangents. | |
| 5: Fully on-topic with no noticeable digressions. | |
| Breadthmeasures how widely a response explores various ideas, concepts, options, or perspectives. | 1: Very narrow; offering no alternative ideas. |
| 2: Covers a small range of ideas, omitting key concepts. | |
| 3: Moderately diverse but limited in scope. | |
| 4: Addresses multiple angles with reasonable detail. | |
| 5: Very wide-ranging; covers a broad spectrum of ideas thoroughly. | |
| Depth evaluates the extent of explanations, expertise, and valuable insights in a response. | 1: Very minimal detail; lacks substance or justification. |
| 2: Superficial treatment with limited insight. | |
| 3: Moderately detailed; some analysis or explanation. | |
| 4: Good explanation with solid reasoning and insight. | |
| 5: Deeper explanation, demonstrated with examples or expertise | |
| Clarity assesses how easily a reader can understand the response, considering conciseness, organization, and readability. | 1: Very unclear; difficult to follow or disorganized. |
| 2: Somewhat unclear; main ideas partially obscured. | |
| 3: Moderately clear; understandable with some effort. | |
| 4: Clear, well-structured, and easy to read. | |
| 5: Flows seamlessly with minimal effort to comprehend. | |
| Actionability assesses whether the response offers guidance readily implementable in a visualization (N/A for guideline-only questions). | 1: No actionable guidance or next steps. |
| 2: Little actionable guidance; mostly explanations. | |
| 3: Some actionable guidance; partially implementable. | |
| 4: Clear, actionable guidance that can be followed easily. | |
| 5: Highly actionable; provides explicit step-by-step instructions. |
The researchers initially assigned scores to the 20% sample, engaging in discussions to resolve any discrepancies and establish consensus definitions and criteria for each score within every metric to ensure consistency. Following this, one of the researchers proceeded to score the remaining question responses. Throughout this coding process, they held continuous meetings to discuss any new issues and questions to ensure consistency and quality of codes and scores, refining metric definitions and score applicability.
Given that some questions elicited multiple Human responses while others received only one, we made them to be evaluated with a single score to assess the collective quality of the Human responses. This score captures the cumulative strengths of all sub-responses, recognizing that a single high-quality sub-response can adequately address the questioner’s needs even if others are less informative. We performed Mann–Whitney U tests to compare Human composite scores with ChatGPT response quality scores.
Furthermore, any elements of the Human responses that extended beyond the capabilities of ChatGPT, such as providing citations to related literature, links to external resources, visual examples, or posing follow-up questions, were recorded. We also categorized whether each question sought feedback on the questioner’s visualizations or inquired about general visualization design knowledge.
@@ 3.3. Analysis Decisions
We selected ChatGPT as a representative of LLM-based chatbots for several reasons. While these chatbots use the same transformer architectures, ChatGPT is the most widely adopted in practice. Also, in our early pilot experimentation with Bard (now Gemini), Claude, and Bing (which uses the same GPT model as ChatGPT), ChatGPT generally provided responses that were better aligned in terms of style and content. For example, Bing often veered off by drawing information from various web sources, which were frequently inferior to ChatGPT’s internal knowledge responses. Furthermore, Bard’s recommendations were narrower in scope and often not applicable, like suggesting the addition of a trend line to a Gantt Chart. Overall, our impression was that these systems are mostly comparable, with ChatGPT having a slight edge over the others. Several studies also indicated that ChatGPT demonstrates superior performance compared to other LLMs in terms of a variety of task questions involving reasoning, facts, and fairness (Borji and Mohammadian, 2023; Ahmed et al., 2023). That said, we discuss an automatic evaluation pipeline, outlined in the discussion section, to enable comparisons of different models without requiring human intervention.
@@
Why use The Six Quality Metrics?
Our goal was to assess the semantic quality of responses to visualization design questions—how well they conveyed meaningful, useful, and understandable knowledge—via human evaluation. While quantifiable measures (e.g., sentence structure, rhetorical style) can provide quick and repeatable approximations of text quality (Van der Lee et al., 2021), research has shown that they do not fully capture the complex and composite nature of constructs like diversity by quantifiable metrics (Van Miltenburg et al., 2018) and do not often correlate with how humans assess text quality (Scott and Moore, 2006; Reiter, 2018). Based on a review of literature on the human evaluation of automatically generated texts, we selected metrics that cover distinct aspects of intrinsic quality. Coverage and topicality capture the nuanced aspects of relevance—how the response answers the user’s questions directly (Van der Lee et al., 2021; Sreedhar et al., 2024; Awasthi et al., 2023); breadth and depth reflect the richness of ideas and reasoning (Wang et al., 2023b); actionability assesses practical applicability and informativeness (Van der Lee et al., 2021); and clarity is a widely used measure to assess readability, organization, and conciseness (Jiang et al., 2024; Awasthi et al., 2023). Overall, this framework offers a meaningful foundation for evaluating LLM performance in visualization design, with future work potentially extending it to include response tone, prioritization, or interaction quality.
@@ 3.4.1. Ratings on Quality Metrics
ChatGPT 4 achieved the highest score of 4.97 ( = 0.22), closely followed by ChatGPT 3.5 with 4.92 ( = 0.33). Humans scored lower at 4.37 ( = 0.95). Mann–Whitney U tests revealed that the difference between both ChatGPT models and Humans was statistically significant (), while the comparison between ChatGPT 3.5 and ChatGPT 4 was not significant ().
The lower score often stemmed from the Human respondent’s failure to address all facets of the question. For instance, in Figure 3C, the user raised a two-part question, “…would you recommend all the states be placed on the Y-axis instead of grouped as regions? If so, which visual layouts would be effective with dealing with the large number of states?…”. However, the Human response only addressed the second part of the question, suggesting a heat map as a useful visual layout.
@@
ChatGPT 4 achieved the highest score of 4.95 ( = 0.31), closely followed by ChatGPT 3.5 with 4.87 ( = 0.45). Humans scored slightly lower at 4.72 ( = 0.60). Mann–Whitney U tests revealed that ChatGPT 4 significantly outperformed Humans (), and ChatGPT 3.5 also showed a statistically significant advantage over Humans (). However, the comparison between ChatGPT 3.5 and ChatGPT 4 was not significant ().
While all conditions performed well in addressing the correct question, there were moments when both ChatGPT and Humans strayed from the main topic. For instance, when a user inquired about the application of visual variables in VR visualizations,111visguides:visual-variables-in-vr ChatGPT 3.5 not only addressed this question but also came up with a hypothetical question asking what factors to consider when choosing such visual variables. Human respondents also occasionally veered off-topic by offering unsolicited advice and extending their responses beyond the original question’s scope.
@@
ChatGPT 3.5 achieved the highest score of 4.62 ( = 0.61), slightly outperforming ChatGPT 4, which scored 4.45 ( = 0.66). Humans scored significantly lower at 3.35 ( = 1.13). Mann–Whitney U tests revealed that both ChatGPT 3.5 and ChatGPT 4 significantly outperformed Humans ( for both comparisons). Additionally, ChatGPT 3.5 demonstrated a statistically significant advantage over ChatGPT 4 ().
For example, in Figure 3A, when the user inquired about color palette guidelines, ChatGPT provided an extensive list of potential approaches, encompassing data type, color harmony, accessibility, etc. In contrast, the Human respondent to this question concentrated primarily on the creation of domain-specific color maps, offering in-depth insights into the process.
Depth
For the depth metric, ChatGPT 4 achieved the highest score of 4.43 ( = 0.56), outperforming both ChatGPT 3.5, which scored 3.44 ( = 0.71), and Humans, who scored 3.52 ( = 1.07). Mann–Whitney U tests revealed that ChatGPT 4 significantly outperformed both ChatGPT 3.5 () and Humans (). However, there was no significant difference between ChatGPT 3.5 and Humans ().
There were intriguing and consistent behavioral patterns across the conditions. For example, ChatGPT 3.5 responses tended to provide general, surface-level improvements applicable to any visualization, such as adding labels and legends (Figure 3). While these suggestions were often hit or miss, they were effective when the questions sought generic advice. In contrast, ChatGPT 4 exhibited a more human-like behavior, offering focused suggestions in a more authoritative manner, showcasing expertise rather than merely enumerating a list of ideas, as illustrated in Figures 3 B and C.
Good Human responses were similar to ChatGPT 4, often even more nuanced, as demonstrated by the recommendation to “put a label before the numerical attribute like ‘Estimated Production: 1000’ (Figure 3B).” This detailed guidance directly pertains to the unique aspects of the visualization in question. However, Human responses also displayed greater variability. For instance, when asked about color clarity and alternative data representation to reduce clutter, a Human response simply confirmed the current design choices are fine222visguides:latitude-distribution-of-solar-and-wind-farms-uk. The lower scores of such poor quality single responses pulled the average down.
@@
Notably, ChatGPT consistently presented its responses in a well-structured list, which is a clear advantage of machine-generated responses. This contributed to ChatGPT 4 achieving the highest clarity score, averaging 4.97 ( = 0.20), closely followed by ChatGPT 3.5 with an average score of 4.93 ( = 0.28). Human responses also demonstrated good clarity, with an average score of 4.30 ( = 0.80). However, Humans received lower ratings when the responses were overly brief, making it difficult to discern specific reference points. Mann–Whitney U tests revealed that both ChatGPT models significantly outperformed Humans (), while the difference between ChatGPT 4 and ChatGPT 3.5 was not statistically significant ().
@@
ChatGPT 4 achieved the highest score, averaging 4.92 ( = 0.38), followed by ChatGPT 3.5 with an average score of 3.99 ( = 0.60). Human responses had the lowest average score of 3.85 ( = 0.94). Overall, ChatGPT 4 performed exceptionally well in providing actionable recommendations, as its responses consistently included focused, practical advice. By contrast, ChatGPT 3.5 often provided broader, less specific suggestions, which reduced its scores due to the generic or vague nature of the advice (Figures 3 B and C), although the list of recommendations often made sense (Figure 3A). The actionability scores for Human responses tend to be low when the users refrain from offering recommendations and instead validate all of the user’s choices333visguides:comparison-of-acute-hepatitis-from-tycho-dataset. Mann–Whitney U tests confirmed that ChatGPT 4 significantly outperformed both ChatGPT 3.5 () and Humans (). However, there was no significant difference between ChatGPT 3.5 and Human responses ().